OpenStudy (anonymous):

How would you find the exact value of tan pi/8

OpenStudy (anonymous):

so tan(3.14/8?

OpenStudy (anonymous):

in previous lessons i learned that you can use sum or difference identities... for ex: cos15= cos(45-30) and yes @bubbles98

OpenStudy (anonymous):

there's the half-angle identity as well

OpenStudy (anonymous):

pi/8 is the same thing as 15 degrees

OpenStudy (anonymous):

so tan(45-30)

OpenStudy (anonymous):

so is what i did is i punch it in the calculator and gor 0.006850524

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

and use difference formula for tan

OpenStudy (anonymous):

@bubbles98 that's not the exact answer

OpenStudy (anonymous):

http://www.clarku.edu/~djoyce/trig/tangentformulas.jpg We'll need the third formula

OpenStudy (anonymous):

of course not just a little step so far

OpenStudy (anonymous):

i thought pi/8 equals 22.5 degrees.

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

you'd have to use the half angle formula then

OpenStudy (anonymous):

http://www.freemathhelp.com/images/halfangles.png 3rd one

OpenStudy (anonymous):

pi/8 is the same thing as (pi/4)/2 right?

OpenStudy (anonymous):

so tan pi/8/2

OpenStudy (anonymous):

so... tan(45/2) = sin(45)/(1+cos(45)

OpenStudy (anonymous):

no, it'd be tan(pi/4/2)

OpenStudy (michele_laino):

please keep in mind that: \[1=\tan(\pi/4)=\frac{ 2 \tan(\pi/8) }{ 1-(\tan(\pi/8))^{2} }\]

OpenStudy (michele_laino):

I used this identity: \[\tan(2x)=\frac{ 2 \tan(x) }{ 1-(\tan(x))^{2} }\]